# Operator Spreading and the Absence of Many-Body Localization

### v2

@@ IV Schrieffer-Wolff transformations
Figure 15: Left: The number of terms in the Hamiltonian with support on $l$ sites after performing the Schrieffer-Wolff transformation (IV.9) up to order $k$ for the XXX chain with $D=5$ and $N=20$ lattice sites. For large $k$, this number increases approximately exponentially with $l$. Right: The corresponding sum of $1$-norms of terms with support on $l$ sites averaged over $30$ realizations. While each term is expected to have an amplitude $\sim D^{-2l}$ if there are no resonances, the exponentially increasing number of terms means that the total coupling of distant sites is no longer controlled by a small parameter. For example, we have $s_{1}^{(1)}(k)\approx s_{7}^{(1)}(k)$ for $k=4$.
@@
One way to obtain these charges is to try to perturbatively construct them in the limit of strong disorder [40, 41, 43]. In this limit, one can think of these charges as the local operators $\sigma^{z}_{i}$ smeared out by the hopping and dressed by the interactions. As in our previous considerations, there is then again an important difference between the non-interacting and the interacting case: While in the former the one-particle operators $\sigma^{z}_{i}=c^{\dagger}_{i}c_{i}-1$ remain one-particle operators, they become many-particle operators in the latter case. As we will see, the perturbative construction of local conserved charges will lead us again to multiple commutators which are closely related to the ones studied in the previous sections.
@@ IV.1 Non-interacting case
First, we consider again the non-interacting case. An ansatz for the operator $T$ can be obtained by commuting $H_{0}$ and $V$. This leads to $T=\sum_{j}A_{j}(S^{+}_{j}S^{-}_{j+1}-S^{+}_{j+1}S^{-}_{j})$. The amplitudes $A_{j}$ can then be determined from Eq. (IV.2) leading to
@@
We note, furthermore, that $\overline{|h_{j+1}-h_{j}|}=2/3$, i.e., the factor $\bar{J}^{(0)}_{j,j+1}/|h_{j}-h_{j+1}|$ is small on average if the disorder strength $D$ is large. However, if two neighboring sites have roughly the same potential, then this factor can become large or might even diverge. Rare resonances thus might destroy the perturbative construction if they proliferate. We will come back to this point below.
@@
In the non-interacting case, it is also well understood that distance resonances have a vanishingly small probability to proliferate [1, 3, 4]. In our construction, this scaling argument arises as follows: Two lattice sites $j$ and $j+\ell$ a distance $\ell$ apart are coupled by an effective long-range hopping $\bar{J}^{(\ell-1)}_{j,j+\ell}(c_{j}^{\dagger}c_{j+\ell}+h.c.)$ with an effective amplitude obtained by $\ell-1$ successive Schrieffer-Wolff transformations. The amplitude of this long-range effective hopping term is given by $\bar{J}^{\ell-1}\sim D^{-2\ell+2}\times f[\{(h^{\ell-1}_{j}-h^{\ell-1}_{k})^{-1}\}]$ where $f$ is some function of the inverse differences of the random fields within the cluster of $\ell$ sites. Now the smallest pairwise difference on this cluster on average scales as $\ell^{-2}$. Thus the effective long-range hopping amplitude scales as $\bar{J}^{(\ell-1)}_{j,j+\ell}\sim D^{-2\ell}\ell^{2}$ and is thus typically exponentially suppressed. Therefore resonances have a vanishingly small probability to proliferate and to destroy the perturbative construction above. This is also supported by our symbolic calculations on finite chains of length $N$. We do successive lowest order Schrieffer-Wolff transformations and stop the transformations if the total 1-norm of the off-diagonal part is below some threshold value $\varepsilon$. We have $\sim N$ possible off-diagonal terms of a given range and if the amplitudes of all those terms, on average, are suppressed as $D^{-2k}$ at order $k$ of the transformation then our condition to stop is $ND^{-2k}<\varepsilon$. We thus expect that we need $k\sim\ln(N/\varepsilon)$ many transformation steps. This is consistent with our observations.
@@ IV.2 Interacting case
The related and most fundamental problem, however, is that lattice sites $j$ and $j+\ell$ a distance $\ell$ apart will be coupled by $\sim 3^{2}4^{\ell-1}$ distinct operators because on each site we can put in principle one of the operators $\{\sigma^{x},\sigma^{y},\sigma^{z},\mathbb{I}\}$ but the string of operators of length $\ell+1$ has to start and end with one of the three Pauli matrices. While not all of these combinations of operators will be realized in a microscopic model such as the XXX chain, there will be exponentially many distinct operators which couple these two sites, see Fig. 15. Thus there is, in general, no reason to expect this perturbative construction to work because the overall effective amplitude of terms coupling these two sites will be $\bar{J}_{j,j+\ell}\sim D^{-\ell}4^{\ell}\sim\mathcal{O}(1)$. I.e., there is no small parameter anymore. This reasoning is supported by the results of symbolic calculations shown in Fig. 15. Here we have calculated the sum $s_{\ell}^{(1)}(k)$ of the amplitudes of all terms obtained in the Schrieffer-Wolff transformation (IV.9) to order $k$ which have support on $\ell$ sites, i.e., which connect sites a distance $\ell-1$ apart. For a periodic chain with even length $N$, there are $\sim N$ distinct pairs of sites which are a distance $21,\cdots,N/$ 2 apart. I.e., for a specific pair the overall coupling is on average $\sim s_{\ell}^{(1)}(k)/N$ and, as can be seen in Fig. 15, is in general not exponentially decreasing with $\ell$. We note that the system considered in Fig. 15 has only $N=20$ sites so that only results up to $\ell\sim 10$ are relevant. That the number of terms generated in such a transformation with support on $\ell$ sites is exponentially increasing with $\ell$ makes it very questionable that a construction of local conserved charges in this perturbative manner is a controlled approximation. This is a major issue even before thinking about the problem of resonances and thermal inclusions (regions for a given disorder configuration which, by chance, are only weakly disordered). It seems to us that the exponential increase of the number of terms in such transformations is a point which has not been properly addressed in the proof of MBL in Refs. [40, 41] where the author seems to be only concerned about the effect of resonances. While it is not impossible that for disorder configurations without resonances and thermal inclusions the perturbative construction of local charges is working in the thermodynamic limit—for example, because of interference effects between different terms coupling the same sites—any proof needs to provide rigorous arguments that the exponentially many terms do not overwhelm the exponential small parameter in the expansion. Such arguments seem to be missing so far. We note that for a finite system the construction will always eventually converge because the finite length provides a cutoff and the norm $s_{\ell}^{(1)}(k)$ will eventually drop off exponentially with $\ell$ as can also be seen in Fig. 15.
@@
The destabilizing effects of resonances and thermal inclusions have received considerable attention in the literature already [50, 51, 35]. From the perspective of perturbative Schrieffer-Wolff transformations, they lead to the following additional problems: The renormalized couplings for an operator acting on $\ell$ lattice sites will in its denominator typically contain a linear combination of all the random fields on these $\ell$ sites. Given that there are exponentially many such operators, the simple scaling argument for the non-proliferation of resonances in the Anderson case is clearly no longer applicable in the interacting case. In particular, thermal inclusions—regions of the lattice which have similar potentials—lead to a further destabilization of the perturbative transformation. This can be seen, for example, in symbolic calculations of the unitary transformation (IV.1) with the operator (IV.9) to higher orders. In configurations where such regions exist, the transformation generates long-range terms with large amplitudes even if the disorder $D$ is extremely strong. If, on the other hand, we exclude configurations with such regions and with any pairwise resonances by hand, then the amplitudes of any individual operator typically decay with the length of its support. Another manifestation of this issue can be seen when performing $k$ successive lowest order Schrieffer-Wolff transformations for a finite chain of length $N$. We want to stop the transformations again once the 1-norm of the off-diagonal part is below some threshold $\varepsilon$. In contrast to the non-interacting case, we now have exponentially many terms connecting sites a given distance apart. If we assume that individual amplitudes in the off-diagonal part are suppressed as $D^{-2k}$, then the condition for stopping the transformations becomes $\text{e}^{N}D^{-2k}<\varepsilon$ and we would thus expect that we need $k\sim N$ steps for $N$ sufficiently large. This is, however, not what we observe for typical disorder configurations. The results of our symbolic calculations point instead to an increase of the number of transformations $k\sim\text{e}^{N}$. This implies that some amplitudes in the off-diagonal part are $\sim\mathcal{O}(1)$ and that we have to eliminate almost all of the off-diagonal terms to achieve convergence.
@@
In summary, the iterative construction of local conserved charges cannot simply be transferred from the non-interacting to the interacting case. There are two main issues: Even if one considers only a lowest order transformation, one is not dealing with a constant number of operators but rather with an exponentially increasing number of operators, both in the Schrieffer-Wolff transformation $T^{(k)}$ as well as in the transformed Hamiltonian $H^{(k)}$. Even if one tries to only perform a single unitary transformation (IV.1) with the operator $T$, the number of distinct operators in the transformed Hamiltonian is increasing exponentially with the support of the operator which thus can overwhelm the exponential decay of the amplitude of individual operators with the length of their support. We note that different but related arguments for the failure of a perturbative construction of conserved charges in the interacting case have recently been made in Ref. [52]. Furthermore, regions with similar lattice potentials result in individual operators with large support which, nevertheless, have large amplitudes even for very strong disorder. The latter issue is related to the well-known avalanche instability [50, 51, 53, 35].
@@ V Conclusions
We have considered operator growth in one-dimensional lattice models and its connection to localization from two different perspectives. First, we studied the multiple commutator $[H,A]^{(k)}$, which appears in the Euclidean time evolution, for a local operator $A$ and a Hamiltonian $H$ which is a sum of nearest-neighbor terms. For a generic system, we have shown that the number of distinct terms generated is exponentially increasing with the order $k$ of the commutator. This leads to a strict bound for the commutator norm $s(k)$ which grows faster than exponential with $k$, a result which was first demonstrated in Ref. [30]. Since this result is based on a simple counting of connected clusters generated by the commutator, it is expected that this bound is asymptotically tight for a generic system. We then provided strict bounds for two special cases. First, we considered the Hamiltonian $H$ of a free-fermion system. In this case, all the terms generated by the commutator of the Hamiltonian with a one-body operator $A$ will also always be one-body operators. For a short-range Hamiltonian, the spatial range of operators generated at order $k$ will be proportional to $k$. The number of distinct terms in the commutator can therefore only grow $\sim k^{2}$ as compared to the exponential growth in the general, interacting case. We also proved an exponential bound for the commutator norm in this case. Second, we considered a localized system, either interacting or non-interacting, and showed that an exponential bound for $s(k)$ exists in this case as well. Physically, this can be understood as follows: a localized system has a set of local conserved charges and the Hamiltonian can be written in terms of these charges. Commuting the Hamiltonian with a local operator $A$ then only involves those charges localized near the site the operator $A$ is acting on. Thus, up to exponential tails, only operators with support on a number of sites which is less than the localization length will significantly contribute to the norm.
@@
Having checked our results for non-interacting cases, we then turned to the Heisenberg (XXX) chain with quenched disorder. This model has been suggested to have a transition from an ergodic to a many-body localized phase at some finite disorder strength $D$. In particular, for the MBL phase an effective Hamiltonian written in terms of localized conserved charges is supposed to exist as in the Anderson case but with couplings between these charges which are exponentially decaying with distance. If true, then our proof for localized phases applies. Thus $s(k)$ can grow at most exponentially in such a phase and $s_{l}(k)\sim\exp(-l)$ at fixed $k$. We used symbolic calculations of the commutator and were able to reach order $k=16$ at which point we are considering terms located on a lattice of $332k+1=$ 33 sites. This is substantially larger than what can be done using exact diagonalizations. We found that the growth of the norm of the commutator $s(k)$ appears to be faster than exponential and consistent with the generic bound for non-localized systems for all considered disorder strengths up to $D=25$. Similarly, we find no indication that $s_{l}(k)\sim\exp(-l)$. To the contrary, we find that $s_{l}(k)/s_{1}(k)$ has a maximum which appears to shift to larger and larger $l$ with increasing $k$. At the very least, we claim that our results clearly demonstrate that there is no convincing case for an MBL phase for system sizes up to $33$ sites and disorder strengths up to $D=25$.
@@
While we have concentrated here on the Heisenberg chain with quenched disorder, we note that in the attempted proof of MBL by Imbrie [40, 41] the transverse Ising chain with random couplings and fields in considered instead. For the latter model, it has recently been rigorously proven [48] that the norm of a local operator when commuted $k$-times with the Hamiltonian grows faster than exponential with $k$ and, in fact, asymptotically like the general bound (II.12). Here, we have shown that if a local operator remains quasi-local when commuted with a Hamiltonian then the total norm cannot grow faster than exponential. If we negate this statement, then this means that if the commutator norm of a local operator with a Hamiltonian grows faster than exponential — as in the transverse Ising case — then the commutator is necessarily not a quasi-local operator. It can at best be algebraically localized. I.e., our results come close to proving that an MBL phase does not exist in the transverse Ising chain. The only assumption we have made is that for the commutator to remain quasi-local, contributions to the connected clusters coming from sites far away from the local operator are exponentially suppressed, see App. A. This seems to be a physical requirement for a localized model. Mathematically it is also hard to see how a local operator can otherwise remain quasi-local. This would then require some miraculous cancellation between different paths which has to happen for any local operator and most disorder configurations. Making these arguments even more rigorous might be the last remaining step in proving that localization is not possible in the many-body case.

### v3

@@ IV Schrieffer-Wolff transformations
Figure 15: Top left: Number of terms in the Hamiltonian with support on $l$ sites after performing the Schrieffer-Wolff transformation (IV.9) up to order $k$ for the XXX chain with disorder and $N=100$ lattice sites. For large $k$, this number increases approximately exponentially with $l$. Note that $\tilde{n}_{l}(k)$ does not depend on the disorder realization; the results are exact. Bottom left: Total norm of terms with support on one site after $k=3$ transformations for $200$ samples versus $\text{max}_{i}|h_{i}-h_{i+1}|^{-1}$. Right: Total norm of terms with support on $l$ sites after $k$ transformations averaged over $200$ realizations for $N=100$ and $D=5$. While each term is expected to have an amplitude $\sim D^{-2l}$ if there are no resonances, the exponentially increasing number of terms means that the total coupling of distant sites is no longer controlled by a small parameter. For example, we have $\tilde{s}_{1}^{(1)}(k)\approx\tilde{s}_{8}^{(1)}(k)$ for $k=4$.
@@
One way to obtain these charges is to try to perturbatively construct them in the limit of strong disorder [47, 48, 50]. In this limit, one can think of these charges as the local operators $\sigma^{z}_{i}$ smeared out by the hopping and dressed by the interactions. As in our previous considerations, there is then again an important difference between the non-interacting and the interacting case: While in the former the one-particle operators $\sigma^{z}_{i}=c^{\dagger}_{i}c_{i}-1$ remain one-particle operators, they become many-particle operators in the latter case. As we will see, the perturbative transformation of the Hamiltonian will lead us again to multiple commutators which are closely related to the ones studied in the previous sections.
@@ IV.1 Non-interacting case
First, we consider again the non-interacting case. An ansatz for the operator $T$ can be obtained by commuting $H_{0}$ and $V$. This leads to $T=\sum_{j}A_{j}(\sigma^{+}_{j}\sigma^{-}_{j+1}-\sigma^{+}_{j+1}\sigma^{-}_{j})$. The amplitudes $A_{j}$ can then be determined from Eq. (IV.2) leading to
@@
We note, furthermore, that $\overline{|h_{j+1}-h_{j}|}=2/3$, i.e., the factor $\bar{J}^{(0)}_{j,j+1}/|h_{j}-h_{j+1}|$ is small on average if the disorder strength $D$ is large. However, if two neighboring sites have roughly the same potential, then this factor can become large or might even diverge. Rare resonances thus might destroy the perturbative construction if they proliferate. An example how such resonances can affect the transformed Hamiltonian in shown in Fig. 15, bottom left panel. We will come back to this point below.
@@
In the non-interacting case, it is also well understood that distance resonances have a vanishingly small probability to proliferate [1, 3, 4]. In our construction, this scaling argument arises as follows: Two lattice sites $j$ and $j+\ell$ a distance $\ell$ apart are coupled by an effective long-range hopping $\bar{J}^{(\ell-1)}_{j,j+\ell}(c_{j}^{\dagger}c_{j+\ell}+h.c.)$ with an effective amplitude obtained by $\ell-1$ successive Schrieffer-Wolff transformations. The amplitude of this long-range effective hopping term is given by $\bar{J}^{\ell-1}\sim D^{-2\ell+2}\times f[\{(h^{\ell-1}_{j}-h^{\ell-1}_{k})^{-1}\}]$ where $f$ is some function of the inverse differences of the random fields within the cluster of $\ell$ sites. Now the smallest pairwise difference on this cluster on average scales as $\ell^{-2}$. Thus the effective long-range hopping amplitude scales as $\bar{J}^{(\ell-1)}_{j,j+\ell}\sim D^{-2\ell}\ell^{2}$ and is thus typically exponentially suppressed. Therefore resonances have a vanishingly small probability to proliferate and to destroy the perturbative construction above. This is also supported by our symbolic calculations on finite chains of length $N$. First, we note that the size of the norm of terms with support on one site is entirely determined by how close to a nearest-neighbor resonance a given sample is. Distant resonances play no role; all the data fall onto a ’hockey stick’ curve, see Fig. 15, bottom left panel. Second, we consider successive lowest order Schrieffer-Wolff transformations and stop the transformations if the total 1-norm of the off-diagonal part is below some threshold value $\varepsilon$. We have $\sim N$ possible off-diagonal terms of a given range and if the amplitudes of all those terms, on average, are suppressed as $D^{-2k}$ at order $k$ of the transformation then our condition to stop is $ND^{-2k}<\varepsilon$. We thus expect that we need $k\sim\ln(N/\varepsilon)$ many transformation steps. This is consistent with our observations.
@@ IV.2 Interacting case
The related and most fundamental problem, however, is that lattice sites $j$ and $j+\ell$ a distance $\ell$ apart will be coupled by $\sim 3^{2}4^{\ell-1}$ distinct operators because on each site we can put in principle one of the operators $\{\sigma^{x},\sigma^{y},\sigma^{z},\mathbb{I}\}$ but the string of operators of length $\ell+1$ has to start and end with one of the three Pauli matrices. While not all of these combinations of operators will be realized in a microscopic model such as the XXX chain, there will be exponentially many distinct operators which couple these two sites, see Fig. 15. Thus there is, in general, no reason to expect this perturbative construction to work because the overall effective amplitude of terms coupling these two sites will be $\bar{J}_{j,j+\ell}\sim D^{-\ell}4^{\ell}\sim\mathcal{O}(1)$. I.e., there is no small parameter anymore. This reasoning is supported by the results of symbolic calculations shown in Fig. 15. Here we have calculated the sum $\tilde{s}_{\ell}^{(1)}(k)$ of the amplitudes of all terms obtained in the Schrieffer-Wolff transformation (IV.9) to order $k$ which have support on $\ell$ sites, i.e., which connect sites a distance $\ell-1$ apart. For a periodic chain with even length $N$, there are $\sim N$ distinct pairs of sites which are a distance $21,\cdots,N/$ 2 apart. I.e., for a specific pair the overall coupling is on average $\sim\tilde{s}_{\ell}^{(1)}(k)/N$ and, as can be seen in Fig. 15, is in general not exponentially decreasing with $\ell$. That the number of terms generated in such a transformation with support on $\ell$ sites is exponentially increasing with $\ell$ makes it very questionable that a construction of local conserved charges in this perturbative manner is a controlled approximation. This is a major issue even before thinking about the problem of resonances and thermal inclusions (regions for a given disorder configuration which, by chance, are only weakly disordered). It seems to us that the exponential increase of the number of terms in such transformations is a point which has not been properly addressed in the proof of MBL in Refs. [47, 48] where the author seems to be only concerned about the effect of resonances. While it is not impossible that for disorder configurations without resonances and thermal inclusions the perturbative construction of local charges is working in the thermodynamic limit—for example, because of interference effects between different terms coupling the same sites—any proof needs to provide rigorous arguments that the exponentially many terms do not overwhelm the exponential small parameter in the expansion. Such arguments seem to be missing so far. We note that for a finite system the construction will always eventually converge because the finite length provides a cutoff and $\tilde{s}_{\ell}^{(1)}(k)$ will eventually drop off exponentially with $\ell$.
@@
The destabilizing effects of resonances and thermal inclusions have received considerable attention in the literature already [56, 57, 42]. From the perspective of perturbative Schrieffer-Wolff transformations, they lead to the following additional problems: The renormalized couplings for an operator acting on $\ell$ lattice sites will in its denominator typically contain a linear combination of all the random fields on these $\ell$ sites. Given that there are exponentially many such operators, the simple scaling argument for the non-proliferation of resonances in the Anderson case is clearly no longer applicable in the interacting case. This can be seen in Fig. 15, bottom left panel, where the sum of the amplitudes of terms in the Hamiltonian with support on one site, $\tilde{s}_{1}(k)$, in the interacting case is large even for samples which do not have nearest-neighbor resonances. Thermal inclusions—regions of the lattice which have similar potentials—lead, furthermore to an additional destabilization of the perturbative transformation. This can be seen, for example, in symbolic calculations of the unitary transformation (IV.1) with the operator (IV.9) to higher orders. In configurations where such regions exist, the transformation generates long-range terms with large amplitudes even if the disorder $D$ is extremely strong. If, on the other hand, we exclude configurations with such regions and with any pairwise resonances by hand, then the amplitudes of any individual operator typically decay with the length of its support. Another manifestation of this issue can be seen when performing $k$ successive lowest order Schrieffer-Wolff transformations for a finite chain of length $N$. We want to stop the transformations again once the 1-norm of the off-diagonal part is below some threshold $\varepsilon$. In contrast to the non-interacting case, we now have exponentially many terms connecting sites a given distance apart. If we assume that individual amplitudes in the off-diagonal part are suppressed as $D^{-2k}$, then the condition for stopping the transformations becomes $\text{e}^{N}D^{-2k}<\varepsilon$ and we would thus expect that we need $k\sim N$ steps for $N$ sufficiently large. This is, however, not what we observe for typical disorder configurations. The results of our symbolic calculations point instead to an increase of the number of transformations $k\sim\text{e}^{N}$. This implies that some amplitudes in the off-diagonal part are $\sim\mathcal{O}(1)$ and that we have to eliminate almost all of the off-diagonal terms to achieve convergence.
@@
In summary, the iterative transformation of the Hamiltonian cannot simply be transferred from the non-interacting to the interacting case. There are two main issues: Even if one considers only a lowest order transformation, one is not dealing with a constant number of operators but rather with an exponentially increasing number of operators, both in the Schrieffer-Wolff transformation $T^{(k)}$ as well as in the transformed Hamiltonian $H^{(k)}$. Even if one tries to only perform a single unitary transformation (IV.1) with the operator $T$, the number of distinct operators in the transformed Hamiltonian is increasing exponentially with the support of the operator which thus can overwhelm the exponential decay of the amplitude of individual operators with the length of their support. We note that different but related arguments for the failure of a perturbative construction of conserved charges in the interacting case have recently been made in Ref. [58]. Furthermore, regions with similar lattice potentials result in individual operators with large support which, nevertheless, have large amplitudes even for very strong disorder. The latter issue is related to the well-known avalanche instability [56, 57, 59, 42].
@@ V Conclusions
We have considered operator growth in one-dimensional lattice models and its connection to localization from two different perspectives. First, we studied the multiple commutator $[H,A]^{(k)}$, which appears in the Euclidean time evolution, for a local operator $A$ and a Hamiltonian $H$ which is a sum of nearest-neighbor terms. For a generic, ergodic system, we have shown that the number of distinct terms generated is exponentially increasing with the order $k$ of the commutator. This leads to a strict bound for the commutator norm $s(k)$ which grows faster than exponential with $k$, a result which was first demonstrated in Ref. [30]. Since this result is based on a simple counting of connected clusters generated by the commutator, it is expected that this bound is asymptotically tight for an ergodic system. This operator growth hypothesis is challenged by results for systems such as the random transverse Ising chain which are supposed to have a localized phase but, nevertheless, always show almost factorial operator growth. Here we systematically addressed this issue. We started by providing strict bounds for two special cases. First, we considered the Hamiltonian $H$ of a free-fermion system. In this case, all the terms generated by the commutator of the Hamiltonian with a one-body operator $A$ will also always be one-body operators. For a short-range Hamiltonian, the spatial range of operators generated at order $k$ will be proportional to $k$. The number of distinct terms in the commutator can therefore only grow $\sim k^{2}$ as compared to the exponential growth in the general, interacting case. We also proved an exponential bound for the commutator norm in this case. Second, we considered a localized system, either interacting or non-interacting, and showed that an exponential bound for $s(k)$ exists in this case as well. Physically, this can be understood as follows: a localized system has a set of local conserved charges and the Hamiltonian can be written in terms of these charges. Commuting the Hamiltonian with a local operator $A$ then only involves those charges localized near the site the operator $A$ is acting on. Thus, up to exponential tails, only operators with support on a number of sites which is less than the localization length will significantly contribute to the norm. We note that an almost factorial versus an exponential growth of the commutator norm also leads to distinctly different Lieb-Robinson bounds and asymptotic behavior of matrix elements in the energy eigenbasis. This is further discussed in App. B and App. C.
@@
Having checked our results for non-interacting cases, we then turned to the Heisenberg (XXX) chain with quenched disorder. This model has been suggested to have a transition from an ergodic to a many-body localized phase at some finite disorder strength $D$. In particular, for the MBL phase an effective Hamiltonian written in terms of localized conserved charges is supposed to exist as in the Anderson case but with couplings between these charges which are exponentially decaying with distance. If true, then our proof for localized phases applies. Thus $s(k)$ can grow at most exponentially in such a phase and $s_{l}(k)\sim\exp(-l)$ at fixed $k$. We used symbolic calculations of the commutator and were able to reach order $k=16$ at which point we are considering terms located on a lattice of $332k+1=$ 33 sites. This is substantially larger than what can be done using exact diagonalizations. We found that the growth of the norm of the commutator $s(k)$ appears to be faster than exponential and consistent with the generic, ergodic bound for non-localized systems for all considered disorder strengths up to $D=25$. Similarly, we find no indication that $s_{l}(k)\sim\exp(-l)$. To the contrary, we find that $s_{l}(k)/s_{1}(k)$ has a maximum which appears to shift to larger and larger $l$ with increasing $k$. At the very least, we claim that our results clearly demonstrate that there is no convincing case for an MBL phase for system sizes up to $33$ sites and disorder strengths up to $D=25$.
@@
While we have concentrated here on the Heisenberg chain with quenched disorder, we note that in the attempted proof of MBL by Imbrie [47, 48] the transverse Ising chain with random couplings and fields in considered instead. For the latter model, it has recently been rigorously proven [37] that the norm of a local operator when commuted $k$-times with the Hamiltonian grows faster than exponential with $k$ and, in fact, asymptotically like the general bound (II.12). Here, we have shown that if a local operator remains quasi-local when commuted with a Hamiltonian then the total norm cannot grow faster than exponential. If we negate this statement, then this means that if the commutator norm of a local operator with a Hamiltonian grows faster than exponential — as in the transverse Ising case — then the commutator is necessarily not a quasi-local operator. It can at best be algebraically localized. I.e., our results come close to proving that the operator growth hypothesis is valid in the strict sense that an almost factorial growth indicates an ergodic systems and that an MBL phase does not exist in the transverse Ising chain. The only assumption we have made is that for the commutator to remain quasi-local, contributions to the connected clusters coming from sites far away from the local operator are exponentially suppressed, see App. A. This seems to be a physical requirement for a localized model. Mathematically it is also hard to see how a local operator can otherwise remain quasi-local. This would then require some miraculous cancellation between different paths which has to happen for any local operator and most disorder configurations. For the transverse Ising chain we already know for sure that such destructive interference effects between different paths do not exist because a gauge can be chosen such that all the terms in the commutator are positive [37]. Making these arguments even more rigorous might be the last remaining step in proving that localization is not possible in general in the many-body case.